Westerville Corp. has collected the following data concerning its maintenance costs for the past six months:
Units Produced Total Cost
July 17,300 $31,620
August 32,900 48,000
September 36,100 55,100
October 22,400 39,000
November 40,300 65,200
December 38,200 62,300

(a1)

Calculate the variable cost per unit using the high-low method. (Round variable cost per unit to 2 decimal places, e.g. 12.25.)
Variable cost $ per unit
Answer #1

Answer

Variable cost per unit by using high low method

The following is the formula of high low method

Y = a + bx

a = fixed cost

Y = total cost

x = number of units

b = variable cost per unit

The following is the formula of Variable cost per unit (b)

b = (cost at highest level of activity - cost at lowest level of activity) / highest level of activity - lowest level of activity

Cost at highest level of activity = $65,200

Cost at lowest level of activity = $31,620

Highest level of activity = 40,300 units

Lowest level of activity = 17,300 units

b = (65,200 - 31,620) / (40,300 - 17,300)

b = 33,580 / 23,000 = 1.46

Variable cost per unit is $1.46

If YOU are like the answer please upvote.